# On the Geometry of Semantics in Next-token Prediction

### v1

@@ Appendix B From Sparsity Language Pattern to Concepts Geometry
Recall the centered data-sparsity matrix $\widetilde{\bm{S}}$ and its SVD
@@
and the singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ are ordered:
@@
Adopting terminology from Saxe et al. [2019], we interpret the columns $\bm{u}_{k}\in\mathbb{R}^{V},{\bm{v}}_{k}\in\mathbb{R}^{m},k\in[r]$ of $\bm{U},{\bm{V}}$ as word and context analyzer vectors for concept $k$. For each word $z\in\mathcal{V}$ and each word-concept $k\in[r]$, the component $\bm{u}_{k}[z]$ represents how present or absent is a word $z$ in context $k$. Respectively for contexts.
Specifically, we think of column dimensions of $\bm{U}$ as semantic dimensions that capture semantic categories.
@@
$\bm{u}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{W}}{\bm{W}}^{\top}\bm{u}_{k}\,$
${\bm{v}}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{H}}{\bm{H}}{\bm{v}}_{k}\,.$
Let’s now simplify these by using the known SVD representation of ${\bm{W}}$ and ${\bm{H}}$. Using this representation (i.e. use ${\bm{W}}\leftarrow{\bm{W}}^{\rm{mm}}$, ${\bm{H}}\leftarrow{\bm{H}}^{\rm{mm}}$) we compute
$\mathds{P}_{\bm{W}}=\mathbf{R}\mathbf{R}^{\top}=\mathds{P}_{\bm{H}}.$
Thus,
@@
Figure 10: Left: Context embeddings; Right: Word embeddings. In both plots, concept representations (blue, green) are computed using Eq. (4). Projections onto these concept axes (${\bm{v}}^{\rm{d}}_{1}$, ${\bm{v}}^{\rm{d}}_{2}$) quantify semantic relevance. The first concept (${\bm{v}}^{\rm{d}}_{1}$, blue line) defines a spectrum from plant-related to animal-related elements, with animals (“hunts in groups”, “dog”) projecting positively and plants (“uses sunlight for food”, “apple”) projecting negatively along the same axis. The second concept (${\bm{v}}^{\rm{d}}_{2}$, green line) represents a continuum from non-mammalian to mammalian traits, with mammalian words and features (“cat”, “stays warm-blooded”) having positive projections and non-mammalian ones (“trout”, “has feathers”) having negative projections. Please note that the concepts are not orthogonal in the plots because it is PCA projection to 2D space.
Appendix C On the hierarchical structure of language
As discussed in Sec. 4.3, language semantics exhibit a hierarchical structure, with more comprehensive semantics involving numerous concepts being inherently more detailed. Each concept’s significance is indicated by its associated singular values; higher values denote concepts critical for differentiating words or contexts. Moreover, as discussed in Sec. 4.4, broader semantic categories are acquired more rapidly, indicating a sequential learning progression from general to more specific concepts. Fig. 11 below provides an illustration of this hierarchical structure.
Figure 11: Hierarchical semantic structure revealed by orthant-Based clustering on Simplified TinyStories. Top: Hierarchical semantic structure emerges as we combine concepts. Each row represents clusters formed by combining an increasing number of concept dimensions (from 1 to 5), with each word cloud displaying the most typical words for a specific sign configuration. Semantic categories become progressively more fine-grained moving downward as additional concept dimensions refine the classifications. Bottom: Detailed view of verb semantics. The combination of the first two concept dimensions ($\mathcal{K}=\{1,2\}$ with configuration $C_{\mathcal{K}}=\{+1,+1\}$) identifies verbs. Adding the third dimension further distinguishes between past tense ($C_{\mathcal{K}}=\{+1,+1,+1\}$) and present tense ($C_{\mathcal{K}}=\{+1,+1,-1\}$). The present tense category is further refined by the fifth dimension to differentiate present action verbs from present continuous (-ing) forms. This hierarchical organization mirrors linguistic grammar structures and supports our hypothesis that language models learn fundamental grammatical distinctions (verb vs. non-verb) before more specific ones (tense and aspect distinctions).
Appendix D Connections to closed-form training dynamics Saxe et al. [2013]
@@ Appendix C On the hierarchical structure of language
The neural-collapse literature has extensively studied Eq. (5) for one-hot $\widetilde{\bm{s}}$, primarily in the balanced case (e.g., Mixon et al. [2020], Han et al. [2021], Súkeník et al. [2023], Tirer and Bruna [2022]) but recently also for imbalanced data (e.g., Liu [2024], Hong and Ling [2023]). Most works focus on global minima of regularized UFM, with less attention to unregularized cases or training dynamics. While some landscape analyses provide partial answers about global convergence (e.g., ), they are limited to regularized cases and don’t characterize dynamics. Even for square loss, where Mixon et al. [2022], Han et al. [2021] study training dynamics—notably Han et al. [2021]’s analysis of the ’central path’ in balanced one-hot cases—these results rely on approximations. Thus, a significant gap remains in understanding UFM training dynamics, even for simple balanced one-hot data with square loss.
By interpreting the UFM with square loss in Eq. (5) as a two-layer linear network with orthogonal inputs, we identify an unexplored connection to Saxe et al. [2013], Gidel et al. [2019]’s analysis. They provide explicit characterization of gradient descent dynamics (with small initialization) for square-loss UFM. The key insight is rewriting (5) as $\sum_{j\in[m]}|\widetilde{\bm{s}}_{j}-{\bm{W}}{\bm{H}}\bm{e}_{j}|^{2}$ with orthogonal inputs $\bm{e}_{j}\in\mathbb{R}^{m}$. This enables direct application of their result, originally stated in Saxe et al. [2013] and formalized in Gidel et al. [2019]. For completeness, we state this here in our setting and terminology as a corollary below.
Proposition 1.
Consider gradient flow (GF) dynamics for minimizing the square-loss NTP-UFM (5). Recall the SVD $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$. Assume weight initialization
@@
This aligns with Zhao et al. [2024]’s regularization-path analysis of NTP-UFM with CE loss, where normalized quantities converge as ${\lambda}\rightarrow 0$ (recall Section 3.2):
@@
Figure 12: (Top) Support matrix and SVD factors of centered support matrix for a synthetic example. (Bottom) Training dynamics of GD minimization of NTP-UFM with square loss (Eq. (5) for two initializations: (i) SVD: initialize ${\bm{W}}$ and ${\bm{H}}$ as per Thm. 1 for $\delta=8$. (ii) Rand: intialize ${\bm{W}}$ and ${\bm{H}}$ random Gaussian scaled to match the norm of SVD initialization. Dynamics with the two initialization are shown in red (SVD) and blue (Rand), respectively. Qualitatively the behavior is similar. Left: Training loss and norms of paramteres. Middle-Left: Convergence of word and context gram-matrices and of logits to the theory predicted by Thm. 1. Middle-Right: Convergence of singular values of logit matrix to those of $\bm{\Sigma}$ (see Thm. 1. Right: Projection of logits to subspace orthogonal to $\bm{U}$ and ${\bm{V}}$; Logits with Rand initialization initially have non-zero projection but it becomes zero as training progresses.
D.1 An example of controlling the rate of learning via reweighting
Consider minimizing the following weighted version of (5):
$\min_{{\bm{W}},{\bm{H}}}~{}~{}\left\|\left(\widetilde{\bm{S}}-{\bm{W}}{\bm{H}}\right)\bm{\Omega}\right\|^{2}\,,$ (10)
where $\bm{\Omega}=\operatorname{diag}([\omega_{1},\ldots,\omega_{m}])$ is a diagonal matrix of weights, one for each context. Here, we consider the STEP-imbalanced one-hot classification setting described in Sec. A.4. Concretely, let the support matrix be
$\bm{S}=\begin{bmatrix}\mathds{I}_{V/2}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}_{V/2\times V/2}\\ \mathbf{0}_{V/2\times RV/2}&\mathds{I}_{V/2}\end{bmatrix}$ (11)
where $R$ is the imbalance ratio and without loss of generality we assumed that the first $V/2$ classes are majorities and that minorities have $1$ example (in our language: context) each. (Thus, the total number of examples is $m=RV/2+V/2=(R+1)V/2$.) Recall that $\widetilde{\bm{S}}=(\mathds{I}_{V}-\frac{1}{V}\mathds{1}_{V}\mathds{1}_{V}^{\top})\bm{S}$.
Recall from Sec. A.4 that $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$ where the singular values and the left singular vectors are explicitly given by:
$\bm{\Sigma}=\operatorname{diag}(\begin{bmatrix}\sqrt{R}\mathds{1}_{V/2-1}&\sqrt{(R+1)/2}&\mathds{1}_{V/2-1}\end{bmatrix}),\qquad\bm{U}=\left[\begin{array}[]{ccc}\mathbb{F}&-\sqrt{\frac{1}{V}}\mathds{1}&\mathbf{0}\\ \mathbf{0}&\sqrt{\frac{1}{V}}\mathds{1}&\mathbb{F}\end{array}\right]\in\mathbb{R}^{V\times(V-1)}\,,$
with $\mathbb{F}\in\mathbb{R}^{V/2\times(V/2-1)}$ an orthonormal basis of the subspace orthogonal to $\mathds{1}_{V/2}$. According to [Thrampoulidis et al., 2022, Lem. A.3] we also have
${\bm{V}}^{\top}=\left[\begin{array}[]{ccc}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]\in\mathbb{R}^{(V-1)\times m}\,.$
We now set the weight matrix as follows:
[…]
Direct calculation yields
${\bm{V}}^{\top}\bm{\Omega}$ $=\left[\begin{array}[]{ccc}\sqrt{\frac{1}{R}}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{2}{R(R+1)V}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{2}{(R+1)V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]$ (16)
$=\operatorname{diag}(\begin{bmatrix}\sqrt{\frac{1}{R}}\mathds{1}_{V/2-1}^{\top}&\sqrt{\frac{2}{R+1}}&\mathds{1}_{V/2-1}^{\top}\end{bmatrix})\underbrace{\left[\begin{array}[]{ccc}\mathbb{F}^{\top}\otimes\mathds{1}_{R}^{\top}&\mathbf{0}\\ -\sqrt{\frac{1}{RV}}\mathds{1}_{RV/2}^{\top}&\sqrt{\frac{1}{V}}\mathds{1}_{V/2}^{\top}\\ \mathbf{0}&\mathbb{F}^{\top}\end{array}\right]}_{=:\widetilde{\bm{V}}^{\top}}=\bm{\Sigma}^{-1}\widetilde{\bm{V}}^{\top}\,.$ (20)
where in the last equation, we recognized the diagonal matrix is equal to $\bm{\Sigma}^{-1}$ and called the other matrix $\widetilde{\bm{V}}\in\mathbb{R}^{m\times(V-1)}$. Thus, the effective sparsity matrix $\widetilde{\bm{S}}_{\bm{\Omega}}=:\widetilde{\bm{S}}\bm{\Omega}$ is equal to
$\widetilde{\bm{S}}_{\bm{\Omega}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}\bm{\Omega}=\bm{U}\widetilde{\bm{V}}^{\top}\,.$
It is now easy to check that $\widetilde{\bm{V}}^{\top}\widetilde{\bm{V}}=\mathds{I}_{V-1}$. Thus, the display above is the SVD factorization of $\widetilde{\bm{S}}_{\bm{\Omega}}$. Note that the singular values of this effective sparsity matrix are all equal, unlike the singular values $\bm{\Sigma}$ of the original data-sparsity matrix $\widetilde{\bm{S}}$. With this at hand, we can show analogous to Proposition 1 that the eigenvalues of ${\bm{W}}(t)$ and ${\bm{H}}(t)$ are now all learned at the same time $T=1$. We also confirm this experimentally in Fig. 13.
Figure 13: (Top) Support matrix and SVD factors of centered support matrix for a one-hot STEP-imbalanced example. (Middle) Training dynamics of GD minimization of NTP-UFM with square loss (Eq. (5); same setup and visualizations as Fig. 12. (Bottom) Training dynamics of GD minimization of NTP-UFM with weighted square loss (Eq. (10)) with weights as in Eq. (12). Note that thanks to the weighting all singular factors are now learned simultaneuously.
Appendix E Extended Discussion on Related Works
Word embeddings and semantic analysis in neural probabilistic language models.
The word2vec architecture Mikolov et al. [2013d, b] and its variants, notably GloVe Pennington et al. [2014], represent seminal early neural probabilistic language models. These simple log-bilinear models, trained on large text corpora, revolutionized word embedding learning. As noted in Zhao et al. [2024], NTP-UFM shares structural similarities with these early models, though in both their work and ours, it serves as a tractable abstraction rather than a practical architecture. Our approach differs by learning both context and word embeddings, following modern practice. The foundational work of Levy and Goldberg [2014] connected word2vec’s geometry to matrix factorization of the pointwise mutual information (PMI) matrix—a specialized word co-occurrence matrix. Subsequent works Levy et al. [2015], Turney and Pantel [2010], Baroni and Lenci [2010] empirically demonstrated semantic interpretations of the PMI matrix’s singular factors and principal components. Building on Zhao et al. [2024], which formalizes a modern version of Levy and Goldberg [2014]’s results, our investigation of concepts differs from this classical literature in two key aspects: (a) We study the NTP setting where both context and word embeddings are learned, yielding concepts that relate to both words and contexts; (b) Our data-sparsity matrix differs fundamentally from classical PMI matrices: it is a centered version of the data support matrix (independent of specific next-token probabilities) and has different structural properties—being orthogonal and non-square, unlike the word2vec setting.
Superposition and feature steering.
Our work was partly motivated by recent compelling literature suggesting that embeddings can be decomposed into linear combinations of a finite set of semantic concepts Bricken et al. [2023], Yun et al. [2023], Park et al. [2023]. These insights from mechanistic interpretability have led to practical applications in "feature steering"—where model behavior can be controlled by manipulating concept representations through addition or subtraction Durmus et al. [2024], Konen et al. [2024]. Our analysis complements the mechanistic interpretability approach by providing a systematic framework for understanding how concepts emerge naturally as principal components from training data statistics. Exploring deeper connections between our theoretical framework and the mechanistic interpretability literature remains an intriguing direction for future work. For completeness, we note that Park et al. [2023, 2024] also investigate geometric properties of concept directions, albeit through fundamentally different technical approaches, assumptions, and perspectives.
Saxe et al.’s closed-form dynamics of two-layer linear network training.
Our work draws inspiration from Saxe et al. [2019]. Conceptually, Saxe et al. [2019] uses a two-layer linear neural network as a theoretical proxy to study the emergence of semantic knowledge in human cognition, providing mathematical justifications for phenomena observed in cognitive semantics literature. A key insight from their work is that even a simple two-layer linear network with orthogonal inputs can yield rich and meaningful conclusions about semantic learning. While two-layer neural networks represent perhaps the simplest instances of non-linear learning, their training dynamics generally remain analytically intractable. However, Saxe et al. [2013] (with aspects later formalized in Gidel et al. [2019]) demonstrated that with square loss, orthogonal inputs, and sufficiently small initialization, these dynamics admit exact closed-form solutions. This mathematical characterization underlies their results on semantic information development through singular factors of the network’s input-output correlation matrix. We make a novel connection to this line of work: the UFM fits perfectly within the framework studied by Saxe et al. [2013, 2019]. Specifically, the UFM can be viewed as a linear two-layer network where the input dimension equals the number of input examples (in our case, the number of contexts $m$). This connection is valuable in two directions: First, the UFM—recently popularized through neural collapse literature (see below)—provides perhaps the most natural and practical setting satisfying Saxe et al.’s seemingly restrictive orthogonal input assumptions. Second, this connection allows us to leverage Saxe et al.’s earlier results in the evolving neural collapse literature. Despite these methodological similarities with Saxe et al. [2019], our work differs in motivation and interpretation. We focus specifically on NTP and how semantic and grammatical concepts emerge from natural language data, rather than general cognitive development. Additionally, we primarily focus on CE loss which is typically used in NTP training.
Neural-collapse geometries.
Our results contribute to the recent literature on the neural collapse (NC) phenomenon Papyan et al. [2020a]. Originally observed in one-hot classification training of DNNs, neural collapse describes two key properties of well-trained, sufficiently expressive DNNs: (1) NC: embeddings of examples from the same class collapse to their class mean, and (2) ETF-geometry: class-mean embeddings form a simplex equiangular tight frame (ETF), being equinorm and maximally separated, with classifier weights exhibiting the same structure and aligning with their respective class-mean embeddings. This phenomenon, consistently observed across diverse datasets and architectures, has sparked extensive research interest, generating hundreds of publications a complete review of which is beyond our scope. Instead, we discuss below the most closely-related line of works. One fundamental direction, which forms the basis for many extensions, focuses on explaining NC’s emergence through the UFM. This model abstracts training as joint optimization of last-layer embeddings (unconstrained by architecture) and classifier weights Mixon et al. [2022], Fang et al. [2021]. Multiple influential works have proven NC emergence by analyzing the UFM’s global optima (e.g., Zhu et al. [2021], Garrod and Keating [2024], Jiang et al. [2023]), with extensions to various loss functions beyond cross-entropy, including square loss Zhou et al. [2022a], Han et al. [2021], Tirer and Bruna [2022], Súkeník et al. [2023, 2024] and supervised contrastive loss Graf et al. [2021], Zhou et al. [2022b], Kini et al. . Most early works maintained the original assumptions from Papyan et al. [2020a]: balanced data (equal examples per class) and embedding dimension $d$ exceeding the number of classes $C$. Recent work has explored $d<C$ settings, though often requiring additional assumptions on the loss function Jiang et al. [2023], Liu et al. [2023]. More substantial progress has emerged in the $d>C$ regime with unbalanced data, where Thrampoulidis et al. [2022] provided a complete characterization for step-imbalanced data (where examples are distributed equally within minority classes and equally within majority classes). They introduced the SELI (simplex-encoding labels interpolation) geometry, showing that logits interpolate a simplex-encoding matrix—a centered version of the one-hot encoding matrix. The embeddings and classifier vectors are then determined, up to rotation and scaling, by the singular vectors of this matrix. The SELI geometry emerges as a special case of the richer geometries characterized in the NTP setting by Zhao et al. [2024]. Together with Li et al. [2023], these works stand alone in extending geometric characterization beyond one-hot encoding—to soft-label and multilabel settings respectively. Specifically, Zhao et al. [2024] analyzes the soft-label setting arising in NTP training on natural language, showing that word and context embeddings are determined by the singular factors of the data sparsity matrix. Our work deepens this understanding by revealing that these SVD factors encode conceptual meaning, thereby extending neural collapse geometry to capture not only the structure of embeddings but also the organization of latent concepts.
Appendix F Additional Experiment Results
F.1 Semantics for Individual Concept

### v2

@@ C.1 Concepts in Embedding Space
This section includes derivations of the formulas of word-concept and context-concept representations of Sec. 3.1. Recall the centered data-sparsity matrix $\widetilde{\bm{S}}$ and its SVD
@@
with singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ that are ordered:
@@
We interpret the columns $\bm{u}_{k}\in\mathbb{R}^{V},{\bm{v}}_{k}\in\mathbb{R}^{m},k\in[r]$ of $\bm{U},{\bm{V}}$ as word and context analyzer vectors for concept $k$. For each word $z\in\mathcal{V}$ and each word-concept $k\in[r]$, the component $\bm{u}_{k}[z]$ represents how present or absent is a word $z$ in context $k$. Respectively for contexts. Thus, we think of column dimensions of $\bm{U}$ as concepts that capture semantic categories. For completeness, we note that each word-analyzer vector $\bm{u}_{k}$ has both positive and negative entries, as it is orthogonal to $\mathds{1}_{V}$ (which lies in the nullspace of $\widetilde{\bm{S}}$).
@@
$\bm{u}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{W}}{\bm{W}}^{\top}\bm{u}_{k}\,\qquad\text{and}\qquad{\bm{v}}^{\rm{d}}_{k}$ $=\mathds{P}_{\bm{H}}{\bm{H}}{\bm{v}}_{k}\,.$
We can further simplify these by using the known SVD representation of $\bm{W}$} and $\bm{H}$} from Sec. 2. Using this representation (i.e. use ${\bm{W}}\leftarrow{\bm{W}}^{\rm{mm}}$, ${\bm{H}}\leftarrow{\bm{H}}^{\rm{mm}}$) we compute $\mathds{P}_{\bm{W}}=\mathbf{R}\mathbf{R}^{\top}=\mathds{P}_{\bm{H}},$ where recall that $\mathbf{R}$ is a partial orthogonal matrix. Thus,
@@
Figure 8: Left: Context embeddings; Right: Word embeddings. In both plots, concept representations (blue, green) are computed using Eq. (5). Projections onto these concept axes (${\bm{v}}^{\rm{d}}_{1}$, ${\bm{v}}^{\rm{d}}_{2}$) quantify semantic relevance. The first concept (${\bm{v}}^{\rm{d}}_{1}$, blue line) defines a spectrum from plant-related to animal-related elements, with animals (“hunts in groups”, “dog”) projecting positively and plants (“uses sunlight for food”, “apple”) projecting negatively along the same axis. The second concept (${\bm{v}}^{\rm{d}}_{2}$, green line) represents a continuum from non-mammalian to mammalian traits, with mammalian words and features (“cat”, “stays warm-blooded”) having positive projections and non-mammalian ones (“trout”, “has feathers”) having negative projections. (The concepts are not orthogonal in the plots because it is PCA projection to 2D space.)
C.2 Proof of Theorem 1
@@
The neural-collapse literature has extensively studied Eq. (6) for one-hot $\widetilde{\bm{s}}$, primarily in the balanced case (e.g., Mixon et al. [2020], Han et al. [2021], Súkeník et al. [2023], Tirer and Bruna [2022]) but recently also for imbalanced data (e.g., Liu [2024], Hong and Ling [2023]). Most works focus on global minima of regularized UFM, with less attention to unregularized cases or training dynamics. While some landscape analyses provide partial answers about global convergence, they are limited to regularized cases and do not characterize dynamics. Even for square loss, where Mixon et al. [2022], Han et al. [2021] study training dynamics—notably Han et al. [2021]’s analysis of the ’central path’ in balanced one-hot cases—these results rely on approximations. Thus, a significant gap remains in understanding UFM training dynamics, even for simple balanced one-hot data with square loss.
By interpreting the UFM with square loss in Eq. (6) as a two-layer linear network with orthogonal inputs, we identify an unexplored connection to Saxe et al. [2013], Gidel et al. [2019]’s analysis. They provide explicit characterization of gradient descent dynamics (with small initialization) for square-loss UFM. The key insight is rewriting (6) as $\sum_{j\in[m]}|\widetilde{\bm{s}}_{j}-{\bm{W}}{\bm{H}}\bm{e}_{j}|^{2}$ with orthogonal inputs $\bm{e}_{j}\in\mathbb{R}^{m}$. This enables application of their result, originally stated in Saxe et al. [2013] and formalized in Gidel et al. [2019]. The following is a detailed (re)statement of Theorem 1.
Theorem 2.
Consider gradient flow (GF) dynamics for minimizing the square-loss NTP-UFM (6). Recall the SVD $\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$. Assume weight initialization (see Fig. 9 for empirical justification of this form of spectral initialization)
@@
This aligns with the regularization-path analysis of NTP-UFM with CE loss, discussed in Sec. 2, where normalized quantities converge as ${\lambda}\rightarrow 0$:
@@
Figure 9: (Top) Support matrix and SVD factors of centered support matrix for a synthetic example. (Bottom) Training dynamics of GD minimization of NTP-UFM with square loss (Eq. (6) for two initializations: (i) SVD: initialize $\bm{W}$} and $\bm{H}$} as per Thm. 2 for $\delta=8$. (ii) Rand: intialize $\bm{W}$} and $\bm{H}$} random Gaussian scaled to match the norm of SVD initialization. Dynamics with the two initialization are shown in red (SVD) and blue (Rand), respectively. Qualitatively the behavior is similar. Left: Training loss and norms of paramteres. Middle-Left: Convergence of word and context gram-matrices and of logits to the theory predicted by Thm. 2. Middle-Right: Convergence of singular values of logit matrix to those of $\bm{\Sigma}$ (see Thm. 2). Right: Projection of logits to subspace orthogonal to $\bm{U}$ and $\bm{V}$}; Logits with Rand initialization initially have non-zero projection but it becomes zero as training progresses.
Appendix D Additional Experimental Results
D.1 Minimal One-Hot Imbalanced Setting
In Sec. 3.3, we identified class imbalance as the minimal requirement for progressive semantic emergence. Here, we provide detailed analysis of this minimal setting.
The emergence of semantically meaningful concepts stems from the structure of labels encoded in $\bm{S}$ (or its centered version $\widetilde{\bm{S}}$) across different contexts. Consider the minimal case: each context has exactly one next-token, with contexts distributed equally across the vocabulary. Here, $\bm{S}$ can be rearranged as $\mathds{I}_{V}\otimes\mathds{1}_{m/V}^{\top}$, yielding trivial concepts where all singular directions contribute equally. This setting corresponds to balanced one-hot classification studied in the neural collapse literature, where last-layer embeddings and weights form highly symmetric aligned structures Papyan et al. [2020a], reflecting the symmetric nature of $\bm{S}$ with no interesting conceptual structure.
However, such balanced settings are unrealistic for language modeling. As seen in the controlled and real experiments in Secs. 3 and 4, natural language produces rich label formations where $\bm{S}$ induces a geometry of concepts that provides semantic interpretation to embedding structures. Here, we demonstrate that meaningful concepts emerge even in one-hot classification—the traditional focus of neural collapse literature—given minimal deviation from perfect balance.
Specifically, consider class imbalance with ratio $R$: $V/2$ majority classes each have $R>1$ times more samples than the $V/2$ minority classes. It can be analytically shown that $\widetilde{\bm{S}}$ has singular values exhibiting a three-tier structure [Thrampoulidis et al., 2022b]:
$\sigma_{1}$ $=\ldots=\sigma_{V/2-1}=\sqrt{R}$ (11)
$>\sigma_{V/2}=\sqrt{(R+1)/2}$ (12)
$>\sigma_{V/2+1}=\ldots=\sigma_{V-1}=1.$ (13)
Moreover, the left singular vectors matrix $\bm{U}$ takes a sparse block form:
where $\mathbb{F}\in\mathbb{R}^{V/2\times(V/2-1)}$ is an orthonormal basis of the subspace orthogonal to $\mathds{1}_{V/2}$. [^4] For concreteness, $\mathbb{F}$ can be constructed using the discrete cosine transform matrix, excluding the constant column: $\mathbb{F}[i,j]=\sqrt{\frac{4}{V}}\cdot\cos\left(\frac{\pi(2i-1)j}{V}\right)$ for $i\in[V/2]$, $j\in[V/2-1]$.
The structure of $\bm{U}$ reveals three distinct concept types, corresponding to the three tiers of singular values:
(1) First V/21V/2-1 columns (largest singular values): Non-zero only for majority classes, representing distinctions among majority classes.
(2) Middle column (singular value $\sqrt{(R+1)/2}$): Opposite-signed entries for majority versus minority classes, encoding the majority-minority dichotomy.
(3) Last V/21V/2-1 columns (smallest singular values): Non-zero only for minority classes, capturing distinctions among minority classes.
This structure demonstrates that the network learns concepts in order of singular value magnitude: first majority class distinctions, then the majority-minority split, and finally minority class differences. Fig. 10 shows an experiment on imbalanced MNIST that confirms this semantic interpretation of concepts (columns of $\bm{U}$) and validates that class imbalance alone suffices for progressive semantic emergence.
Figure 10: Experimental illustration of the fact that important concepts are learned first. Confusion matrix evolution during training of a 3-layer convolutional network on an imbalanced MNIST data with $V=4$ classes, 2 majorities, 2 minorities and imbalance ratio $R=10$. The matrices show five snapshots at different training steps (0, 13, 30, 55, 65) for four classes: two majority classes (Maj1, Maj2) with 100 samples each and two minority classes (Min1, Min2) with 10 samples each. Training progresses from left to right: initially classifying all data as Maj1 (Step 0), then correctly identifying majority classes while misclassifying minority classes (Step 13), gradually improving minority class recognition (Step 30), confusion between minority classes only (Step 55), and finally achieving perfect classification by Step 60. For this setting, Thrampoulidis et al. [2022a] show the singular values of $\widetilde{\bm{S}}$ exhibit a three-tier structure: $\sigma_{1}=\ldots=\sigma_{V/2-1}=\sqrt{R}>\sigma_{V/2}=\sqrt{(R+1)/2}>\sigma_{V/2+1}=\ldots=\sigma_{V}=1.$ Additionally, The structure of $\bm{U}$ reveals three distinct types of concepts, corresponding to the three tiers of singular values: (1) First V/21V/2-1 columns (non-zero only for majority classes) represent distinctions among majority classes; (2) Middle column (singular value $\sqrt{(R+1)/2}$) has opposite-signed entries for majority versus minority classes, encoding the majority-minority dichotomy; (3) Last V/21V/2-1 columns (non-zero only for minority classes) capture distinctions among minority classes. Thus, the evolution of confusion matrices in the figure perfectly aligns with this hierarchical structure.
D.2 Training transformer on Simplified Tinystories dataset
For Simplified TinyStories, we train 6-layer and a 12-layer decoder-only transformers, with 8 attention heads, and 1024 hidden dimensions. Training uses the AdamW optimizer with learning rate 1e-3, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay 0.001. We use batch size 64 and train until convergence (approximately 1000 epochs).
D.3 On the hierarchical structure of language
As discussed in Sec. 3.2, language semantics exhibit a hierarchical structure, with more comprehensive semantics involving numerous concepts being inherently more detailed. Fig. 11 below provides an illustration of this hierarchical structure.
Figure 11: Hierarchical semantic structure revealed by orthant-based clustering on Simplified TinyStories. Top: Clusters are formed by combining 1 to 5 concept dimensions; each row shows word clouds for different sign configurations. Finer-grained semantic categories emerge as more dimensions are added. Bottom: Verb-related structure: $\mathcal{K}={1,2}$ highlights verbs; adding a third dimension separates past and present tense; the fifth dimension further distinguishes verb aspects.
D.4 Semantics for Individual Concept